Solve the following equations:
(i) (D² + DD′ + D + D′ + 1)z = 0
(ii) (D² − 2DD′+ D′²)z = tan(y + x).
1
Expert's answer
2020-03-16T13:18:08-0400

Equation (i)

"\left(D^2+DD'+D+D'+1\right)z=0"

Let us give some theory to understand how to solve this equation:

Put "D=h" and "D'=k" in "f(D, D')=0" , then we get an

auxillary equation is


"f(h, k)=0 \quad\quad\quad\quad(1)"

Solve "(1)" and find the value of "h" interm of k or "k" interm of "h" . Let the "n" values of "h" be denoted by "f_1(k)" , "f_2(k)" , "f_3(k)" , . . . , "f_n(k)" .

Then


"C.F.=\sum C_1e^{f_1(k)x+ky}+\ldots+\sum C_ne^{f_n(k)x+ky}"

In our case,

"\left(D^2+DD'+D+D'+1\right)\equiv h^2+hk+h+k+1=0\\[0.3cm]
h^2+(1+k)h+(k+1)=0-\text{quadratic equation}\\[0.3cm]
\sqrt{D}=\sqrt{(k+1)^2-4(k+1)}=\sqrt{(k+1)(k-3)}\\[0.3cm]
\left[\begin{array}{l}
h_1=\displaystyle\frac{-(k+1)+\sqrt{(k+1)(k-3)}}{2}\\[0.3cm]
h_2=\displaystyle\frac{-(k+1)-\sqrt{(k+1)(k-3)}}{2}
\end{array}\right."

Then,


"C.F.=\sum C_1e^{h_1(k)x+ky}+\sum C_2e^{h_2(k)x+ky}=\\[0.3cm]
=\sum C_1e^{x\cdot\left(\frac{-k-1-\sqrt{(k+1)(k-3)}}{2}\right)+ky}+\sum C_2e^{x\cdot\left(\frac{-k-1+\sqrt{(k+1)(k-3)}}{2}\right)+ky}=\\[0.3cm]
=e^{-x/2}\cdot\sum C_1e^{x\cdot\left(\frac{-k-\sqrt{(k+1)(k-3)}}{2}\right)+ky}+\\[0.3cm]
+e^{-x/2}\cdot\sum C_2e^{x\cdot\left(\frac{-k-1+\sqrt{(k+1)(k-3)}}{2}\right)+ky}=\\[0.3cm]
=e^{-x/2}\cdot\left(f_1\left(L_1x+ky\right)+f_2\left(L_2x+ky\right)\right)"

Conclusion,

"\left(D^2+DD'+D+D'+1\right)z=0\longrightarrow\\[0.3cm]
C.F.=e^{-x/2}\cdot\left(f_1(L_1x+ky)+f_2(L_2x+ky)\right),\text{where}\\[0.3cm]
\left[\begin{array}{l}
L_1=\displaystyle\frac{-k+\sqrt{(k+1)(k-3)}}{2}\\[0.3cm]
L_2=\displaystyle\frac{-k-\sqrt{(k+1)(k-3)}}{2}
\end{array}\right."

Question (ii)

"(D^2-2DD'+D'^2)z=\tan(x+y)\longrightarrow\\[0.3cm]
z=C.F.+P.I."

1 STEP: We try to find C.F.


"(D^2-2DD'+D'^2)z=0"

The auxillary equation is


"m^2-2m+1=0\longrightarrow (m-1)^2=0\longrightarrow\\[0.3cm]
m_{1,2}=1"

Then,


"C.F.=f_1(y+1\cdot x)+x\cdot f_2(y+1\cdot x)\\[0.3cm]
\boxed{C.F.=f_1(y+x)+x\cdot f_2(y+x)}"

2 STEP: We try to find P.I.

Let us give some theory to understand how to solve this equation:

If "F(x,y)" is any function, resolve "F(D, D')" into linear factors say


"(D\u2212m_1D'),(D\u2212m_2D'), . . . ,(D\u2212m_nD')"


then


"P.I.=\frac{1}{(D−m_1D')(D−m_2D'). . .(D−m_nD')} F(x, y)"

Note:

"(1)\,\,\,\frac{1}{(D−mD')}F(x, y)=\int F(x, c−mx)dx,\,\,\text{where}\,\,y=c−mx\\[0.3cm]
(2)\,\,\,\frac{1}{(D+mD')}F(x, y)=\int F(x, c+mx)dx,\,\,\text{where}\,\,y=c+mx"



In our case,


"P.I.=\frac{1}{(D-1\cdot D')^2}\tan(x+y)=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\cdot\frac{1}{(D-1\cdot D')}\tan(x+y)=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\cdot\int\tan(x+c-x)dx=[x+y=c]=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\cdot\int\tan(c)dx=\frac{1}{(D-1\cdot D')}\cdot x\tan(c)=\\[0.3cm]
=\int x\tan(c)dx=\frac{x^2}{2}\tan{c}+C_1\equiv\frac{x^2}{2}\tan(x+y)+C_1"


Conclusion,


"\boxed{P.I.=\frac{x^2}{2}\cdot\tan(x+y)+C_1}"

General conclusion,


"z(x,y)=C.F.+P.I.\longrightarrow\\[0.3cm]
\boxed{z(x,y)=f_1(y+x)+x\cdot f_2(y+x)+\frac{x^2}{2}\cdot\tan(x+y)+C_1}"

ANSWER

Question(i)

"\left(D^2+DD'+D+D'+1\right)z=0\longrightarrow\\[0.3cm]
C.F.=e^{-x/2}\cdot\left(f_1(L_1x+ky)+f_2(L_2x+ky)\right),\text{where}\\[0.3cm]
\left[\begin{array}{l}
L_1=\displaystyle\frac{-k+\sqrt{(k+1)(k-3)}}{2}\\[0.3cm]
L_2=\displaystyle\frac{-k-\sqrt{(k+1)(k-3)}}{2}
\end{array}\right."

Question(ii)

"z(x,y)=f_1(y+x)+x\cdot f_2(y+x)+\frac{x^2}{2}\cdot\tan(x+y)+C_1"